Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 5/9/2017

Location : CR21 at 14:00 pm

Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions
    • Give LH2 documentation to Steve to add to official documentation list (Craig, Josef): completed draft exists but further changes expected; stands
    • Update shift checklist to focus on alarms : (Steve): stands (need to sit down with Paolo)
    • Arrange a geometry survey for September just before running starts (Colin/ChrisR) : Tony Millington generally available, fix date closer to the time; stands
  • DC update
  • Hydrogen system update and plans (JBoehm)
  • Status of alarm clearance (DC/Ajit)
  • Update on tracker readout (Melissa)
  • Update on DS Shield pressure (Henry)
  • Update on SS QPS connector issue (Craig)
  • Cycle 2017/01 Issues
  • Hall/Operations plan : attachment:17-08-17-CW.xlsx
  • Operations startup
    • Initial discussion on system readiness : Wednesday August 23rd in CR25 in R1 @ 11
    • Pre-startup meeting : Monday Sep 4th (Postponed until after the pulser run)
    • Pulser run : Thursday 7th : test of DAQ and online. Paolo will oversee this starting with a phone discussion
      to plan the day on Wednesday @ 16:00
  • Update on Financial Restrictions In Technology and Engineering Subjects (CWhyte)
    No.
  • AOB

Minutes:

Present: MU,KL,HN,AN,CM,CW,Ajit,Paolo,Durga

Actions
Give LH2 SHIFTER documentation to Steve to add to official documentation list (Josef): completed draft exists but further changes expected from Josef & Andy; stands
Arrange a geometry survey for September just before running starts (Colin/ChrisR) : Tony Millington generally available, fix date closer to the time; stands

DC update

Hydrogen system update and plans (JBoehm)
Vent line leak checking in progress
Calibration of H2 sensors to take place this Friday.
Crane stops to go in place asap. Upper (east end) crane needs to be disabled.
Matt Hills now positioned to help out as on-call expert. Shown around on Tuesday.
Safety tour went well with only a few minor outstanding issues. Completed sign-off is expected in the next few days.
When leak checking complete, the aim is too cool with He over the weekend and then start with H2 first thing Monday.
Need magnet and LH2 babysitters next weekend.

Status of alarm clearance (DC/Ajit) - Ajit working through list. Many are known accepted faults. Disabling some others.
Update on tracker readout (Melissa) - No warm up now. MU and PK coming on Thurs for pulser run. Will put plan regarding waveguide swaps together then.
Update on DS Shield pressure (Henry) - Jerg had a look at DS and thinks problem relates to absorber contamination. Proposes flush with N2. Action Thurs/Fri
Update on SS QPS connector issue (Craig) - Pulled out current measurement unit and clicked on SSU connector properly. Seems slightly dodgy though presently stable. Will monitor from here.

Hall/Operations plan : 17-08-17-CW.xlsx

  • Pulser run : Thursday 7th : test of DAQ and online. Paolo will oversee this starting with a phone discussion
    to plan the day on Wednesday @ 16:00

AOB
Hall trench water is 27C, HN says this is understood and ok
New script with Nagios looks at security probe devices.
This Friday LH2 celebration lunch at 12:30pm in R9

Updated by MacWaters, Craig about 5 years ago · 5 revisions